OpenStudy (anonymous):

Can someone help me with this graphing problem Solve and graph x + y = 0 x - y + 2 = 0

OpenStudy (netlopes1):

Please, look this attach file.

OpenStudy (anonymous):

Thank you so much. I have another question how do you graph equations in that format ?

OpenStudy (anonymous):

Is there a formula?

OpenStudy (netlopes1):

no, there isn't a formula, only you must plot (draw) each function. See, for example, "x+y=0" is the same thing that y=-x and "x-y+2=0" is y=x+2, and (-1,1) is the intersection point of two function. Understood? Sorry my english, please.

OpenStudy (anonymous):

oh okay i understand it now

OpenStudy (anonymous):

thank you for explaining

OpenStudy (netlopes1):

Thanks and good studies...